# A Perfectly Truthful Calibration Measure

### v1

@@ Appendix A Non-Truthfulness of Known Calibration Measures
In this section, we prove Theorem A.2 showing that condition (1) holds for a broad family of calibration measures: $\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal}$ and $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$ (Definition A.1), where $\alpha\geq 1$ is arbitrary. Condition (1) states that, for any realization of the states, reporting the average over predictions is weakly better than reporting truthfully for known calibration measures. By our discussion in Section 1.1 after (1), this proves that these calibration measures are not truthful.
@@
Table 1 provides an example illustrating the non-truthfulness of known calibration measures and the truthfulness of our $\mathrm{ATB}$. The table compares two strategies: predicting the average over ground truths and predicting truthfully.
@@
Table 1: The calibration errors of predictors with two samples. The ground truth distribution of the two samples are $25\%$ and $75\%$, respectively. In the table, avg stands for the uninformative predictor that always outputs $50\%$ and truth stands for the truthful predictor that outputs $25\%$ and $75\%$. We calculate the error of the predictors given each realization of the state and the total expected error. For non-truthful error metrics, the expected error of a truthful predictor is strictly higher than the expected error of an uninformative predictor. For $\mathrm{ATB}$, the expected errors are the same.
Appendix B Standard Uniform Convergence Bounds
@@
Definition B.1 (Rademacher complexity).
@@
Proposition B.2 (Uniform convergence from Rademacher complexity).
@@
Proposition B.3 (Rademacher Complexity after Lipschitz Postprocessing).
@@
Definition B.4 (VC Dimension (Vapnik and Chervonenkis, 1971)).
@@
Proposition B.5 (Rademacher Complexity from VC Dimension).

### v2

@@ Appendix A Non-Truthfulness of Known Calibration Measures
In this section, we prove Theorem A.2 showing that condition (1) (restated below as condition (45)) holds for a broad family of calibration measures: $\ell_{\alpha}\textup{-}\textsc{ECE},\ell_{\alpha}\textup{-}\textsc{BinECE},\textsc{smCal}$ and $\ell_{\alpha}\textup{-}\underline{\textsc{distCal}}$ (Definition A.1), where $\alpha\geq 1$ is arbitrary. Condition (1) states that, for any realization of the states, reporting the average over predictions is weakly better than reporting truthfully for known calibration measures. By Remark A.3, this proves that these calibration measures are not truthful.
@@
Remark A.3.
It is very easy to find (many) examples of $\bm{y}$ where the inequality (45) becomes strict, in which case reporting $\bar{\bm{r}}$ instead of the $\bm{r}$ is strictly better (i.e. $\bar{\bm{r}}$ dominates $\bm{r}$). In particular, we can find many examples of $\bm{y}$ where $\textsc{Cal}(\bm{r},\bm{y})>0$ and $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$. This is because for all of the calibration measures Cal mentioned above (including the continuous ones), we always have $\textsc{Cal}(\bm{r},\bm{y})>0$ as long as $r_{1},\ldots,r_{T}$ are distinct values [^5] For binned calibration error, we need $\max_{t}r_{t}-\min_{t}r_{t}$ to be sufficiently large so that $r_{1},\ldots,r_{T}$ do not fall in the same bin. in, say, $[1/3,2/3]$, and we have $\textsc{Cal}(\bar{\bm{r}},\bm{y})=0$ as long as the average outcome $\frac{1}{T}\sum_{t=1}^{T}y_{t}$ is equal to $\bar{r}$.
@@
Table 1 provides an example illustrating the non-truthfulness of known calibration measures and the truthfulness of our $\mathrm{ATB}$. The table compares two strategies: predicting the overall average and predicting truthfully.
@@
Table 1: The calibration errors of predictors with two data points. The ground truth probabilities of the two points are $25\%$ and $75\%$, respectively. In the table, avg stands for the uninformative predictor that always outputs $50\%$ and truth stands for the truthful predictor that outputs $25\%$ and $75\%$. We calculate the error of the predictors given each realization of the state and the total expected error. For non-truthful error metrics, the expected error of a truthful predictor is strictly higher than the expected error of an uninformative predictor. For $\mathrm{ATB}$, the expected errors are the same.
Appendix B Deferred Proofs in Section 5
B.1 Decision-Theoretic Explanation of Blackwell Dominance
In this section, we explain the connection of proper losses and decision making. A proper loss evaluates the loss of a decision maker when they best responds to a prediction.
Definition B.1 (Decision problem).
A decision problem consists of an outcome space $\mathcal{Y}$, an action space $A$, and a loss
$L:\;A\times\mathcal{Y}\to\mathbb{R},\qquad(a,y)\mapsto L(a,y).$
Given a prediction $p\in\Delta(\mathcal{Y})$, the expected loss of an action $a\in A$ under $p$ is
$R_{L}(a;p)\;:=\;\operatorname{\mathbb{E}}_{y\sim p}\!\big[L(a,y)\big].$
A best response to $p$ is any action
$\text{BR}_{L}(p)\;\in\;\operatorname*{arg\,min}_{a\in A}R_{L}(a;p),$
and the best-responding loss, which we will refer to as the induced loss is a function of the prediction and the outcome
$l_{L}(p,y)\;:=\;L\big(\text{BR}_{L}(p),\,y\big),\qquad(p,y)\in\Delta(\mathcal{Y})\times\mathcal{Y}.$
Theorem B.2 shows that the space of induced best-responding loss is the same as the space of proper losses. The proof of this connection can be found in previous work, e.g., Frongillo and Kash (2014); Kleinberg et al. (2023); Hu and Wu (2024).
Theorem B.2 (Induced Losses $\Leftrightarrow$ Proper Losses).
Induced losses and proper losses are equivalent in the following sense:
  • For any decision problem $(A,\mathcal{Y},L)$, the induced loss $l_{L}(p,y)=L(\text{BR}_{L}(p),y)$ is proper.
  • Conversely, for any proper loss $l$, there exists a decision problem $(A,\mathcal{Y},L)$ whose induced loss equals $l$; in particular, one can take $A=\Delta(\mathcal{Y})$ and $L(a,y)=l(a,y)$, so that $l_{L}=l$. If $l$ is strictly proper, then $\text{BR}_{L}(p)=\{p\}$ for all $p$.
B.2 Proof of Theorem 5.2
The proof of Theorem 5.2 relies on a characterization of Blackwell dominance that constructs a coupling between the two predictors, shown in Lemma B.3. The error is comparable under the coupled joint distribution, shown in Lemma B.4.
Lemma B.3 (Blackwell’s Informativeness Theorem).
Assume predictors $A$ and $B$ are both calibrated. Let $J^{A}$ be the joint distribution of $(r^{A},y)\in\Delta(\mathcal{Y})\times\mathcal{Y}$ for the prediction of predictor $A$ and the state. Similarly, $J^{B}$ for predictor $B$. The following three conditions are equivalent:
  1. 1.
    AA Blackwell dominates $B$;
  2. 2.
    There exists a coupling $\Pi$ of the the random variables $r^{A}$, $r^{B}$, and state $y$, such that: $r^{B}$ is conditionally independent of $y$ given $r^{A}$;
  3. 3.
    For the same coupling above, $\operatorname{\mathbb{E}}[r^{A}|r^{B}]=r^{A}$ holds almost surely.
Lemma B.4.
Let Cal be an error measure that is ex-ante truthful (Definition 2.12). Let $\Pi$ be a joint distribution of $(r^{A},r^{B},y)\in[0,1]\times[0,1]\times\{0,1\}$ such that the distribution of $(r^{A},y)$ is calibrated, and $r^{B}$ is independent of $y$ conditioned on $r^{A}$. For $T$ i.i.d. examples $(r^{A}_{1},r^{B}_{1},y_{1}),\ldots,(r^{A}_{T},r^{B}_{T},y_{T})$ drawn from $\Pi$, defining $\bm{r}^{A}:=(r^{A}_{1},\ldots,r^{A}_{T}),\bm{r}^{B}:=(r^{B}_{1},\ldots,r^{B}_{T}),\bm{y}:=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{r}^{A},\bm{y}^{A})]\leq\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{r}^{B},\bm{y}^{B})].$ (53)
Proof of Lemma B.4.
By our assumption that $r^{B}$ is independent of $y$ given $r^{A}$, and that the distribution of $(r^{A},y)$ is calibrated, we have
$\operatorname{\mathbb{E}}[y|r^{A},r^{B}]=\operatorname{\mathbb{E}}[y|r^{A}]=r^{A}$
Define feature $x:=(r^{A},r^{B})$. The equation above becomes $r^{A}(x)=\operatorname{\mathbb{E}}[y|x]$. This means that $r^{A}$ can be viewed as the ground-truth predictor. For $n$ i.i.d. examples $(r^{A}_{1},r^{B}_{1},y_{1}),\ldots,(r^{A}_{n},r^{B}_{n},y_{n})$ drawn from $\Pi$, define $x_{t}:=(r^{A}_{t},r^{B}_{t})$ for $t=1,\ldots,T$ and define $S:=((r^{A}_{t},r^{B}_{t}))_{t\in[T]}$. By the truthfulness of Cal, we have
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{A})]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r^{B})],$
which is equivalent to (53). ∎
Proof of Theorem 5.2.
By Blackwell’s informativeness theorem (Lemma B.3), there exists a coupling between the two predictors such that $r^{\prime}(x)$ is conditionally independent of $y$ given $r(x)$. Thus, the proof is completed by Lemma B.4. ∎
B.3 Proof of Theorem 5.3
Proof of Theorem 5.3.
By Lemma B.4, it suffices to show that $v$ and $y$ are conditionally independent given $\widehat{v}$. Since $y\in\{0,1\}$ is binary, it suffices to show that $\operatorname{\mathbb{E}}[y|v,\widehat{v}$] is determined by $\widehat{v}$ alone and does not additionally depend on $v$, which is indeed true:
$\operatorname{\mathbb{E}}[y|v,\widehat{v}]=\operatorname{\mathbb{E}}[y|v]=\widehat{v}$$\qed$
Appendix C $\mathrm{ATB}$ and Brier Loss
For ATB as a special case of UBSE, we have the following stronger result than Lemma 3.6:
Lemma C.1 (ATB and Brier loss).
Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$. For $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ drawn i.i.d. from $J$, defining $\bm{v}:=(v_{1},\ldots,v_{T}),\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\mathrm{ATB}(\bm{v},\bm{y})]\geq\frac{1}{T}\operatorname{\mathbb{E}}_{(v,y)\sim J}[(v-y)^{2}].$ (54)
The inequality becomes an equality if $J$ is perfectly calibrated.
Proof.
For every fixed threshold $q\in[0,1]$, we have
$\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}<q}(v_{t}-y_{t})\right)^{2}\right]$ $=\operatorname{\mathbb{E}}\left[\left(\sum_{t=1}^{T}(v_{t}-y_{t})\mathbb{I}\left[v_{t}<q\right]\right)^{2}\right]$
$=T\operatorname{\mathbb{E}}_{J}[((v-y)\mathbb{I}\left[v<q\right])^{2}]+T(T-1)\operatorname{\mathbb{E}}_{J}[(v-y)\mathbb{I}\left[v<q\right]]^{2}$ (because $(v_{1},y_{1}),\ldots,(v_{T},y_{T})$ are drawn i.i.d. from $J$)
$\geq T\operatorname{\mathbb{E}}_{J}[((v-y)\mathbb{I}\left[v<q\right])^{2}]$
$=T\operatorname{\mathbb{E}}_{J}[(v-y)^{2}\mathbb{I}\left[v<q\right]].$
Similarly,
$\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}\geq q}(v_{t}-y_{t})\right)^{2}\right]\geq T\operatorname{\mathbb{E}}_{J}[(v-y)^{2}\mathbb{I}\left[v\geq q\right]].$
Summing up the two inequalities above, for every $q\in[0,1]$ we have
$\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\left(\sum_{t:v_{t}<q}(v_{t}-y_{t})\right)^{2}+\left(\sum_{t:v_{t}\geq q}(v_{t}-y_{t})\right)^{2}\right]\geq\frac{1}{T}\operatorname{\mathbb{E}}_{J}[(v-y)^{2}].$
Taking expectation over $q\sim\mathrm{Unif}([0,1])$ proves (54). When $J$ is perfectly calibrated, all inequalities in this proof become equalities, so (54) also holds as an equality. ∎
Appendix D Standard Uniform Convergence Bounds
@@
Definition D.1 (Rademacher complexity).
@@
Proposition D.2 (Uniform convergence from Rademacher complexity).
@@
Proposition D.3 (Rademacher Complexity after Lipschitz Postprocessing).
@@
Definition D.4 (VC Dimension (Vapnik and Chervonenkis, 1971)).
@@
Proposition D.5 (Rademacher Complexity from VC Dimension).